11 points value a. The cost ot a new automobile is $11,500 It the interest rate is 8%, how much would you have to set aside now to provide this sum in eight years? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Present value 6216.20 b. You have to pay $10,000 a year in school fees at the end of each of the next nine years, if the interest rate is 11%, how much do you need to set aside today to cover these bills? (Do not round Intermediate calculations. Round your answer to 2 decimal places.) Present value I С You have invested $90 000 at 11%. A er paying the above sch00 fees, how much would remain at the end ofthe nine years? o no round intermediate calculat ns. Round o rans vero decimal places.) Future value ReferenceseBook & Resources
Answer #1
1)
a. Present Value $       6,213.09
Working:
# 1 Present value of 1 = (1+i)^-n Where,
= (1+0.08)^-8 i 8%
= 0.540269 n 8
# 2 Present Value = $           11,500 * 0.540269
= $       6,213.09
b. Present Value $     63,676.05
Working:
# 1 Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.11)^-9)/0.11 i 11%
= 5.537048 n 9
# 2 Present Value = $           11,500 * 5.537048
= $     63,676.05
c. Future Value $ 2,30,223.32
Working:
# 1 Future Value of 1 = (1+i)^n Where,
= (1+0.11)^9 i 11%
= 2.55804 n 9
# 2 Future Value = $           90,000 * 2.55804
= $ 2,30,223.32